OpenStudy (anonymous):

The French Club is sponsoring a bake sale. If their goal is to raise at least $140, how many pastries must they sell at $3.50 each in order to meet that goal? Write and solve an inequality. A. 140p3.50; p40 B. 3.50p140; p490 C. 3.50p140; p136.5 D. 3.50p140; p40

OpenStudy (anonymous):

40 is how many they need to sell

OpenStudy (anonymous):

so its a or d

OpenStudy (anonymous):

The goal is 140, and they want to make AT LEAST that amount (sounds a little like greater than or equal to), and sell the pastries at 3.50 EACH (sounds like that's where the x would go).

OpenStudy (anonymous):

140 / 3.50 = 40 so 40 is a or d

OpenStudy (anonymous):

\[140 \le 3.50x\]

OpenStudy (anonymous):

so after u know that 140/3.50= ? that leves u with 2 answers which are ?

OpenStudy (anonymous):

@trexnapora

OpenStudy (anonymous):

send me a message when ur there

OpenStudy (anonymous):

Above is the inequality, 40 is the number they need to sell.

OpenStudy (anonymous):

she's not there